OpenStudy (anonymous):

someone guide me through this problem. I dont know how to solve it. [(-1-i)^2 x (2i)^2]/(2-2i)

jimthompson5910 (jim_thompson5910):

I'm guessing they want you to rationalize the denominator. So you will have to multiply top and bottom by 2+2i. This is the conjugate of 2-2i

OpenStudy (anonymous):

Convert all complex numbers to trigonometric form and then simplify the expression. Write the answer in standard form.

jimthompson5910 (jim_thompson5910):

oh gotcha

OpenStudy (anonymous):

I tried to rationalize... I got a huge mess... thats why I am hoping someone can guide me through it

jimthompson5910 (jim_thompson5910):

were you able to convert (-1-i)^2 to trig form?

OpenStudy (anonymous):

well i didnt do that part yet bc of the number next to it

jimthompson5910 (jim_thompson5910):

-1-i is in the form a+bi a = -1 b = -1 theta = arctan(b/a) = arctan(-1/(-1)) = arctan(1) = 45 degrees theta is in Q3, so add 180: 45+180 = 225 degrees

jimthompson5910 (jim_thompson5910):

r = sqrt(a^2 + b^2) r = sqrt((-1)^2 + (-1)^2) r = sqrt(2)

jimthompson5910 (jim_thompson5910):

therefore, -1-i converts to sqrt(2)*[cos(225) + i*sin(225)]

jimthompson5910 (jim_thompson5910):

and you can use DeMoivre's theorem to find [sqrt(2)*[cos(225) + i*sin(225)]]^2

OpenStudy (anonymous):

wait ... why would the entire formula be sqred? it is multiplying something else that needs to be sqred

jimthompson5910 (jim_thompson5910):

well initially (-1-i) is squared so because it is equivalent to sqrt(2)*[cos(225) + i*sin(225)], I'm squaring all of sqrt(2)*[cos(225) + i*sin(225)]

OpenStudy (anonymous):

ok but what abt the 2isqrd

jimthompson5910 (jim_thompson5910):

well (2i)^2 is simply 4i^2 you square 2 to get 4 you square i to get i^2

jimthompson5910 (jim_thompson5910):

but I guess you could also convert that to trig form too

OpenStudy (anonymous):

-4 but then what

jimthompson5910 (jim_thompson5910):

2i = 0+2i a = 0 b = 2 theta = arctan(b/a) = arctan(2/0) = undefined so theta is 90 degrees r = 2 because 2i is 2 units away from the origin

OpenStudy (anonymous):

ok what do i do with that?

OpenStudy (anonymous):

dont know what to do with 0+2i how would i put it into the form... of 2(cos90?

jimthompson5910 (jim_thompson5910):

2i = 2*[cos(90) + i*sin(90)]

jimthompson5910 (jim_thompson5910):

so far we have -1-i = sqrt(2)*[cos(225) + i*sin(225)] and 2i = 2*[cos(90) + i*sin(90)]

jimthompson5910 (jim_thompson5910):

what do you get when you square sqrt(2)*[cos(225) + i*sin(225)]

OpenStudy (anonymous):

2 cos90 + isin90

OpenStudy (anonymous):

so now i multiply them??

OpenStudy (anonymous):

both sets?

jimthompson5910 (jim_thompson5910):

yes but after you square sqrt(2)*[cos(225) + i*sin(225)]

jimthompson5910 (jim_thompson5910):

what did you get for that piece?

OpenStudy (anonymous):

thats what i got... 2cos 90....

OpenStudy (anonymous):

because it went to 450 but then i used the coterminal

jimthompson5910 (jim_thompson5910):

oh gotcha, so (-1-i)^2 = 2*[cos(90) + i*sin(90)] I see what you meant now

jimthompson5910 (jim_thompson5910):

if 2i = 2*[cos(90) + i*sin(90)] then what is (2i)^2 equal to

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

if you squared 2*[cos(90) + i*sin(90)], what do you get

OpenStudy (anonymous):

oooo 4(cos180

jimthompson5910 (jim_thompson5910):

so (2i)^2 = 4*[cos(180) + i*sin(180)]

jimthompson5910 (jim_thompson5910):

now you multiply 2*[cos(90) + i*sin(90)] and 4*[cos(180) + i*sin(180)]

jimthompson5910 (jim_thompson5910):

do you know the shortcut to do so?

OpenStudy (anonymous):

u mean just multiply? 8cis270?

jimthompson5910 (jim_thompson5910):

yeah 2cis(90) with 4cis(180) to get 8cis(270), good

jimthompson5910 (jim_thompson5910):

so all that work transforms each piece in the numerator to trig form, then you multiply to get 8cis(270)

jimthompson5910 (jim_thompson5910):

now convert the denominator 2-2i to cis form

OpenStudy (anonymous):

i got 2sqrt2 cis 45

jimthompson5910 (jim_thompson5910):

2-2i is not in Q1, it's in Q4

jimthompson5910 (jim_thompson5910):

so the angle isn't 45, but it is ???

OpenStudy (anonymous):

2sqrt2cis 135

jimthompson5910 (jim_thompson5910):

well you were on the right track with the 45, but it's really -45 -45 + 360 = 315 so 2-2i = 2*sqrt(2)*cis(315)

OpenStudy (anonymous):

okok got it.. i was confused...

jimthompson5910 (jim_thompson5910):

so up top, things simplified to 8*cis(270) in the denominator, we have 2*sqrt(2)*cis(315)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

i figured it out.. (well mainly thanks to u) thank you so very much

jimthompson5910 (jim_thompson5910):

what answer did you get

OpenStudy (anonymous):

so it came otu to be 2-2i

jimthompson5910 (jim_thompson5910):

yeah all of that and it turns out the answer is buried in plain sight: in the denominator lol

OpenStudy (anonymous):

yes I was thinking that... but its not always the case